OpenStudy (anonymous):

Find the value of two numbers if their sum is 12 and their diffference is 4.

OpenStudy (anonymous):

8 and 4

OpenStudy (anonymous):

Set up two equations and solve. x+y=12 x-y=4 Solve for y in one: y=12-x Plug that into the other equation: x-(12-x)=4 x-12+x=4 2x-12=4 2x=16 x=8 Then plug x into one of the equations you solved for y: y=12-8 y=4 Soooooo, the two numbers are 8 and 4.